A mortar shell is projected at an angle of 65° above the horizontal and then it finally stricken a tower, which is located 25m away. The bullet hits the tower at about 15 m above 5he line of projection. Determine the time taken to reach the tower.


1
Expert's answer
2021-08-25T17:23:29-0400
"v=\frac{25}{t\cos{25}}\\-15=\frac{25}{t\cos{25}}t\sin{25}-0.5(9.8)t^2\\t=2.3\ s"